• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 00:09:40 on March 4,2009

    Entry number 264448

    keyword=Swing shift summary

    Swing shift summary:

    Shift worksers: J. P. Chen (SL), X, Zhan (TO), B. Zhao (3rd)

    16:00 continue production on positive LHRS, LHRS kin F, p=1.13 GeV/c
    Deadtime has been high (30%), people in backroom invistigating.
    BB Sync problem, 3 figures flat line.

    16:20, Stopped BB run per run coordinator's instruction,
    changed PS2 from 1 to 2. new run deadtime ~4%, occationally fluctuate to 60%.
    16:35-16:40 requested beam off for Alexandre to check Compton electron detector.
    17:00-17:25 coda stuck, reboot all components, restarted coda
    (note: when it can not connect to TS1, it is actually TS11).
    18:00 started Moller (Sasha takes over).
    21:40 Moller completed. Beam tuning finished, MCC called.
    Checked BPM positions ok.
    21:45-22:15 Had problems to turn on BB Hvs and VDC HVs. Closed BB HV guis and restarted. BB HVs are fine. Reset crate VDC HVs and they turned back to 4 kV.
    22:20, with 5uA on 120 psig N2, take data on Kin F.
    22:40 moved to pol. He3, changed to kin. D (LHRS=0.9 GeV/c)
    23:45 moved to N2, 5 uA on 120 psig N2.

    BB runs
    1952 15:35-16:20, 4M, deadtime ~30%, 15 uA on Pol. He3.
    1953 16:20-16:33, 0.9M, changed PS2 from 1 to 2, deadtime ~4%
    1954 16:41-17:03, 1.1M, coda problem near end-of-run.
    1955 17:25-17:55, 1.8M.
    1959 22:22-22:34, 1.46M, 5 uA on 120 psig N2.

    1960, 22:41-23:22, 5M. 15 uA on pol. He3.
    1961, 23:23-23:44, 2.2M.
    1962, 23:52- still going on at 0:08. 5 uA on 120 psig N2.

    LHRS runs
    20499, 15:41-16:20, 800K, P=1.13 GeV/c (Kin. F), positive polarity. 15 uA on Pol. He3.
    20500, 16:22-16:33, 262K,
    20501, 16:41-17:03?, no end-of-run, coda problem.
    20502, 17:22-17:55, 720K.
    20505, 22:22-22:34, 480K, 5 uA on 120 psig N2. Kin. F.

    20506, 22:44-23:00, 800K, 15 uA on pol. He3. Kin. D (0.9 GeV/c).
    20507, 23:00-23:17, 800K,
    20508, 23:17-23:36, 800K.
    20509, 23:36-23:44, 330K.
    20510, 23:45-23:46, 5 uA on 120 psig N2. kin D, deadtime 32%.
    20511, 23:47-0:07, 625K, 5 uA on N2, changed PS3 from 1 to 2.